OpenStudy (anonymous):

1. How many moles of glucose (C6H12O6) are in 6.0 liters of a 3.5 M C6H12O6 solution? (3 points) 21 moles 9.5 moles 3.5 moles 1.7 moles Multiple Choice 2. What property of a solution remains the same after it is diluted? (3 points) amount of solvent total volume amount of solute concentration Multiple Choice 3. How many moles of sodium chloride (NaCl) solute are in 155 grams of an 85.5 percent by mass solution? (3 points) 1.19 mol NaCl 2.27 mol NaCl 7.75 mol NaCl 10.6 mol NaCl

OpenStudy (anonymous):

please help :/

OpenStudy (anonymous):

1) 3.5M glucose solution tells that in each Liter there are 3.5 moles of solute. so 3.5*6=21 2)by reducing the concentration/diluting a solution the amount of solvent is increased relative to the solute. 3)%mass NaCl= 155g*0.855=132.525/58.44=2.27 mol NaCl

OpenStudy (anonymous):

thsank you!